OpenStudy (anonymous):

Really weird question, please help! Thank You! A chef makes Chinese Noodles. He first makes a rope of dough(roughly circular cross section, treat as circular) 5.0 cm in diameter. He then doubles the rope on itself (making 2 ropes out of the original rope), repeating (making 4 ropes, etc.). What is the thickness of the original noodle if he folds the noodles 20 times. Thank You! This question has me stumped!

OpenStudy (anonymous):

@triciaal @johnweldon1993 @jabez177 @jim_thompson5910 @SolomonZelman @zepdrix @tkhunny @whpalmer4

OpenStudy (neonumbrella5115):

So let me figure what this question is even asking! It's a bit tricky in wording.  So when the chef doubles the rope, what affect does that have on the diameter of the rope? I'm thinking dividing in half?

jimthompson5910 (jim_thompson5910):

I agree with @neonumbrella5115 the first original rope has a circular cross section with diameter 5 cm then you cut the rope in half, maybe along this line shown here  so you'll now have 2 ropes with circular cross sections of diameter 5/2 = 2.5 cm repeat and you'll go from 2.5 to 2.5/2 = 1.25 etc etc

jimthompson5910 (jim_thompson5910):

You'll notice the diameters follow a geometric sequence, so you can use this formula \[\Large a_n = a*r^{n-1}\] in this case, a = 5 is the original starting diameter and r = 1/2 = 0.5 is reflecting the fact that you cut the diameter in half over and over

OpenStudy (anonymous):

oh I see

OpenStudy (anonymous):

but wouldn't I need to use an equation for cylinders to solve it?

OpenStudy (anonymous):

@neonumbrella5115 yea, I think it divides in 1/2

OpenStudy (whpalmer4):

@jim_thompson5910 I wonder about this answer (and the problem in general). If you have a cylinder with diameter \(d\) and length \(l\), the volume is \[V = \frac{\pi}4 d^2l\]If you then stretch it to make it twice the length with constant volume, \(d\) changes by a factor of \(\sqrt{2}\), not a factor of \(2\). I think this is an example of someone trying to fit a math concept to a real-world item and not doing a good job. The noodles in question are flat, not round. Here's a video showing them being made, note that the process of this problem starts at about the halfway point in the video. https://www.youtube.com/watch?v=JyNN3heu6nM